# Witnessing Disorder in Quantum Magnets

### v2

@@ VI.2 TLL and RS are multipartite entangled
Figure 3: Finite residual tangle (RT) indicating the existence of multi-partite entanglement in the Tomonaga Luttinger liqud (TLL) and random singlet (RS) state. RT of TLL is approaching close to it’s Bethe Ansatz value as N increases. For the RS, RT remains fairly constant.
@@
Figure 4: Observing the growth of entanglement depth at the $q=\pi$ point in Tomonga Luttinger liquid (TLL) and random singlet (RS) state. (a) Quantum Fisher Information (QFI) density for TLL and RS state plotted as function of $q$ for $N=16,20,24,28,30$. (b) QFI density plotted as function of $N$ at $q=\pi$ point for TLL and RS state. The rate of growth of entanglement depth is suppressed in the RS state. Here we observe the TLL state to be at-least $6$-partite entangled while the RS is at least $4$ partite entangled.
To further understand the entanglement structure of these states next, we employ the multi-partite measures, i.e RT and QFI. We find that both the TLL and RS state show a finite multi-partite entanglement as seen through the RT Fig. 3 and QFI Fig. 4.
@@
it’s value in the thermodynamic limit, calculated using Bethe Ansatz [88, 87].
On the other hand, for the RS state, the RT is fairly constant evidencing the existence of multi-partite entanglement. This is contrary to the SDRG picture, where the RS state is pictured as a tensor product of singlets spanning the chain, per disorder realization [63].
Accordingly, a straightforward implication of such a picture is that the RT should vanish for the RS state. However, for the system sizes that we can access, we do not observe such a vanishing behavior Fig. 3.
The multi-partite nature of the TLL [81] and RS is also predicted by the QFI. At the $q=\pi$, we observe that the TLL is at-least 6-partite entangled and that the RS state is at-least 4-partite entangled for the largest spin chain size ($N=30$) we access. From Fig. 4 one can also see an increasing entanglement depth with system size both in the TLL and RS state at the $q=\pi$ point. This growth, however, is more pronounced in the TLL as compared to the RS state. The implication of such a result is that if one observes a multi-partite entanglement using QFI in an experiment, then one cannot simply rule out the RS state as one might have inferred from it’s schematic picture.
In the thermodynamic limit, one expects that the QFI for the TLL to diverge logarithmically [89]. On the other hand, for the RS state, a complete suppression of QFI to a constant value is expected [90, 85].
@@ VII Finite Temperature results
Having discussed the zero temperature entanglement structure of the two states, we move to explore the effects of finite temperature on this structure. For the task at hand, we employ concurrence and the equal-time structure factor as entanglement witness based measures. In Sec. VII.1, we discuss the importance of realizing what is being experimentally accessed and how it effects the way one performs the disorder-average in their analysis. In particular, we highlight the importance of knowing how to perform the disorder-average keeping in mind the observables accessible in an experiment. We point that this is especially crucial when one is trying to rule out disorder induced states by using certain entanglement measures. Following which, in Sec. VII.3, we find that the low-temperature growth of entanglement witness based measure turns out to be very useful to characterize these states.
@@ VII.1 The order of disorder-average matters for concurrence
If one goes by our previous results for concurrence then one expects to see beyond nearest neighbor 2-partite entanglement in the RS state. Does this imply that if experimentally we do not observe further nearest neighbor concurrence can we rule out the presence of a disorder induced state? On the practical side, can the further nearest neighbor entanglement even be observed in INS? Unfortunately, the answer to both these questions turns out to be negative. Typically, in experimental situations such as in INS, we access disorder-averaged spin correlations. What this means is that when characterizing entanglement from spin correlation for example, one does not access the entanglement per disorder realization. While this distinction might feel academic, for measures like concurrence, the order of how one performs to disorder-average becomes important if one is comparing a theoretical model against the experimental predictions.
To begin with, let’s look at the finite temperature behavior of concurrence with the way we have been evaluating it as explained in Sec. VI.1. In Fig. 5, we find that upto some threshold temperatures the RS state has finite pairwise entanglement from further neighbors whereas the TLL state has only nearest neighbor pairwise contribution consistent with what we saw before.
@@
Figure 5: Demonstrating how the order of disorder-average affects the concurrence distribution in the random singlet (RS) state. (a) Disorder-averaged concurrence observed for $1$st, $3$rd and $5$th neighbors as function of temperature for $N=20$. (b) Concurrence, obtained from disorder-averaged spin correlations, observed for the $1$st, $3$rd and $5$th neighbors as a function of temperature. For (a), concurrence is evaluated using Eq. (13) for every disorder realization and then averaged. On the other hand, for (b), concurrence is evaluated using Eq. (47). in which case the pairwise entanglement distribution of the RS state mimics the Tomonga Luttinger liquid (TLL) state i.e., becomes nearest neighbored. To help read the plots some of the error bars coming from disorder averages have been suppressed.
@@
where $\bar{g}^{zz}_{ij}$ is the disorder-averaged spin correlation function evaluated as per Eq. (35). From the experimental perspective, it is this concurrence that would be accessed. In Fig. 6, we perform this and find that RS has only nearest neighbor 2-partite entanglement just like the TLL. We do not observe any finite beyond nearest neighbor concurrence as one might have expected. What this means is that experimentally one would not be able to observe beyond nearest neighbor pairwise entanglement if they access disorder-averaged spin correlations. Thus ruling out the RS state by looking at the 2-tangle or concurrence and observing no beyond nearest neigbor 2-partite entanglement is incorrect. We also observe in Fig. 5b that the nearest neighbor concurrence vanishes at the same temperature for both the TLL and RS state.
Figure 6: Extracting central charge information from the low-temperature behavior of concurrence for the Tomonga Luttinger liquid (TLL) and random singlet (RS) state respectively. The plots here are evaluated for $N=20$ spin chain. For the TLL state, we make use of Eq. (48) and Eq. (49) to fit the observed low-temperature behavior. For the RS state, we make use of Eq. (48) and Eq. (52) for the observed low-temperature behavior. Note that here we are evaluating the concurrence from disorder-averaged spin correlations [Eq. (47)] as one would access in experiments like inelastic Neutron scattering (INS). For ease of viewing, the error bars have been suppressed.
@@ VII.2 The low-temperature behavior of concurrence contains central charge
On the other hand, it has been shown that the RS state is characterized by an effective central charge [92]
@@
We conjecture that it is possible to access this information from the concurrence of the RS state where we employ the disorder-average that is experimentally relevant. We find the following ansatz fits very well with the observed low-temperature behavior for the RS state
@@ VII.3 The low-temperature growth of multi-partite entanglement is a useful probe
We saw that the zero temperature QFI, Fig. 4, predicts both the TLL and RS state to be multi-partite entangled. Thus, stating whether a state is $m$-partite or some other is $l$-partite is not a very useful probe if we want to distinguish them experimentally. Instead, we propose to investigate the low-temperature behavior of multi-partite entanglement witness measures. We employ the equal time structure factor for this task. As evident in Figs. 7, the growth of entanglement at low-temperatures is qualitatively quite different. For the TLL, the low-temperature scaling behavior is well known [93]
@@
On the other hand, we find the growth of entanglement to be quite slow in our disorder induced state. We propose the following emperical behavior,
@@
Figure 7: Comparison of the low-temperature behavior of the equal time structure factor for the Tomonga Luttinger liquid (TLL) and random singlet (RS) state. (a) Equal time structure factor of both the TLL and RS states observed as function of $\log(\beta)$ for $N=16,20,24$ chains. (b) Testing the low-temperature behavior of equal time structure factor for the TLL against the analytic prediction Eq. (55). (c) Testing the low-temperature behavior of equal time structure factor for the RS against the ansatz Eq. (57) for $N=20$ spin chain. The fitting parameters are given in Eq. (56) and Eq. (58). The fits were checked within a fixed interval $1.90.09<\log(\beta)<$ 1.9 for both the TLL and RS state. This choice was made owing to the finite size effects in the TLL state for temperatures below the lower bound. For ease of viewing, the error bars have been suppressed.
@@ VIII Discussion
VIII.1 Qualitative pictures at long length scales can be misleading
We highlight the need to be careful with qualitative pictures when it comes to ruling out RS states. Within the SDRG approach, the RS state is pictured a tensor product of pairwise entangled sites per disorder realization Fig. 8a. Consequently, this approach rules out the existence of any entanglement beyond the pairwise setting. On the other hand, both RT Fig. 3 and QFI Fig. 4, point to the RS state to be multi-partite entangled. In particular, we observed that at $q=\pi$, the QFI density predicts the RS state to be at-least 4-partite entangled for the largest spin chain size $N=30$ that was accessed. Clearly, there is a paradox here.
VIII.2 Multi-partite entanglement localization and understanding RS ‘state’
Figure 8: An illustration of the ground state of the disordered spin chain. (a) The schematic picture of the random singlet (RS) state, as described within the strong disorder renormalization group (SDRG), where pairs of entangled sites are connected by orange colored arcs. (b) Schematic picture of the RS state predicted from our results, where multi-partite clusters (mc) are represented by green shaded ovals of size $l_{mc}$ alongside pairwise entangled sites connected by orange colored arcs. These mc are localized to distances below the crossover length $L_{c}$.
What is the reason for this apparent contradiction? The confusion stems from what one means by the ‘state’ here. Whether one means the critical ground state of the Hamiltonian having information about all the length scales or only the effective state that characterizes the long wavelength physics. These are not the same, especially with regards to entanglement.
The SDRG framework provides an extremely good description of the the long-wavelength behavior of a disordered spin chain [94, 95]. The success of a picture based on a product of singlets implies that the entanglement needed to describe long–wavelength physics is at most 2-partite. However this long–wavelength picture, in which correlations are controlled by the strong–disorder fixed point, should be understood to apply beyond a crossover length, $L_{c}$ [94]. This crossover can be observed in Fig. 4a, as a momentum
@@ VIII.2 Multi-partite entanglement localization and understanding RS ‘state’
above which the RS state and the TLL are practically indistinguishable.
At distances shorter than $L_{c}$, spins may be entangled in clusters and exhibit multipartite entanglement, as illustrated in Fig. 8b. However for distances $\gg L_{c}$, these multi-partite clusters become irrelevant, and entanglement is at most 2-partite. It follows that one should interpret the effective RS state obtained from SDRG as being defined above $L_{c}$. At short length scales, the multi-partite entanglement plays a role, as witnessed by the QFI density [Fig. 4a]. We can understand this as the localization of the multi-partite entanglement.
We can turn this observation into a bound on QFI density. If the RS state, as understood within the SDRG, is at most 2-partite entangled, this places a limit on the entanglement contributed by long lengthscales. We therefore expect
@@
A violation of this inequality would rule out the RS state.
To summarize the above discussion the ground state of a disordered spin chain, the ‘true’ RS state is a multi-partite entangled state. While the multi-partite entanglement is not relevant to understand the behavior it does exist. This multi-partiteness is resultant from contributions below the crossover length i.e., from the short range physics.
VIII.3 Infinite disorder and Thermodynamic limit.
@@
Figure 9: Observing the decreasing entanglement depth with increasing $\delta$ in the random singlet (RS) state. (a) Quantum Fisher information (QFI) density as function of q for $\delta=1.0,1.5,2.0,2.5$ in a $N=30$ spin chain. (b) Residual tangle (RT) as a function of $\frac{1}{\delta}$ for $N=30$ chain, where the line is used as a guide. While an increasing $\delta$ reduces the entanglement depth, for $\delta=2.5$ and $N=30$ spin chain the QFI density at the $q=\pi$ still detects a multi-partite entangled RS state.
@@
where $C^{*}$ is some constant. For real materials, a finite $\delta$ would already imply an entanglement depth larger than two. For instance, in [96], it was reported that Ba5CuIr3O12 single crystals can characterize a RS state with a $\delta\approx 2.5$.
VIII.4 Experimentally ruling out RS states is a subtle business
@@
The RS state, by which we mean the ground state of the disordered spin chain, is not a simple a tensor product of singlets. This could have been inferred from the results in [90, 97], in hindsight. While the 2-partite entanglement of this state may indeed be the relevant entanglement setting to understand the effects of disorder, it does also have a multi-partite contribution from length scales below the crossover length. Even though the multi-partite nature of the state may not be relevant to understand the effects of disorder at long length scales, these contributions do not simply disappear either. Consequently, ruling out RS states under the assumption that it has only 2-partite entanglement is incorrect. Furthermore, we showed how one might be misled into ruling out RS state by the absence of any concurrence beyond the nearest neighbors. One might incorrectly conclude the absence of long range singlets from such an observation.
@@
Convex roof measures like concurrence depend quite a lot on the way the disorder-average is performed. This is due to the presence of the ‘$\max$’ function in the definition Eq. (10). The operation of averaging over disorder realizations does not commute with the max operation. Whether one takes the disorder-average inside the max function, or outside it, the end results are completely different. Why is this important? For quenched disordered systems, the spin correlations are already disorder-averaged. Thus if one were to access spin correlations from experiments like INS and from those tried evaluated concurrence using Eq. (47) then one would observe Fig. 5b as opposed to Fig. 5a. The former corresponds to the case when concurrence has been evaluated from disorder-averaged spin correlations and the latter is the one where we evaluate concurrence for each disorder realization and then average it. In the former case, the RS state shows no further neighbor concurrence, essentially mimicking the TLL state at the 2-partite level Fig. 2c. In [57], the authors incorrectly ruled out RS states by observing vanishing concurrence and multi-partite entanglement using QFI. However, both of these observations cannot rule out the RS state as shown in Sec. VI.2 and VII.1. This point will be discussed further in a follow-up work [98].
VIII.5 What can be useful experimentally?

### v3

@@ VI.2 TLL and RS are multipartite entangled
Figure 3: Finite residual tangle (RT) indicating the existence of multi-partite entanglement in the Tomonaga Luttinger liquid (TLL) and random singlet (RS) phase. RT of TLL is approaching close to its Bethe Ansatz value as N increases. For the RS, RT remains fairly constant.
@@
Figure 4: Observing the growth of entanglement depth at the $q=\pi$ point in Tomonga Luttinger liquid (TLL) and random singlet (RS) phase. (a) Quantum Fisher Information (QFI) density for TLL and RS phase plotted as function of $q$ for $N=16,20,24,28,30$. (b) QFI density plotted as function of $N$ at $q=\pi$ point for TLL and RS phase. The rate of growth of entanglement depth is suppressed in the RS phase. Here we observe the TLL phase to be at-least $6$-partite entangled while the RS is at least $4$ partite entangled.
To further understand the entanglement structure of these phases next, we employ the multi-partite measures, i.e RT and QFI. We find that both the TLL and RS phase show a finite multi-partite entanglement as seen through the RT Fig. 3 and QFI Fig. 4.
@@
it’s value in the thermodynamic limit, calculated using Bethe Ansatz [95, 93]. On the other hand, for the RS phase, the RT is fairly constant evidencing the existence of multi-partite entanglement. The multi-partite nature of the TLL [85] and RS is also predicted by the QFI. At the $q=\pi$, we observe that the TLL is at-least 6-partite entangled and that the RS phase is at-least 4-partite entangled for the largest spin chain size ($N=30$) we access. From Fig. 4 one can also see an increasing entanglement depth with system size both in the TLL and RS phase at the $q=\pi$ point. This growth, however, is more pronounced in the TLL as compared to the RS phase.
The implication of such a result is that if one observes a multi-partite entanglement using QFI in an experiment, then one cannot simply rule out the RS phase as one might have prematurely inferred from it’s schematic picture. This is not in contradiction with the SDRG method. The fixed point of the SDRG method, the infinite randomness fixed point (IRFP), is a description of dominant physics at long length scales, whereas the multipartite entanglement observed here pertains to short length scales (discussed further in Sec VIII). The specific characteristics of this multi-partite entanglement will in general depend on the disorder distribution. Nevertheless, this multi-partite information is encoded within the ground state of the disordered spin chain and has consquences when trying ruling out RS phase experimentally. Whether there are further corrections to the structure of the RS state [75] at longer length scales is something we cannot answer from the system sizes we investigate.
In the thermodynamic limit, one expects that the QFI for the TLL to diverge logarithmically [96]. On the other hand, for the RS phase, a complete suppression of QFI to a constant value is expected [94, 89].
As a historical footnote, we remark that the presence of multipartite entanglement in the RS state could have been anticipated from the arguments of Fisher, who envisaged random singlets forming between effective spin–1/2 objects of “finite extent” [65]. However at the time the SDRG method was introduced, ideas of entanglement yet to permiate beyond the quantum information community.
@@ VII Finite Temperature results
Having discussed the zero temperature entanglement structure of the two phases, we move to explore the effects of finite temperature on this structure. For the task at hand, we employ concurrence and the equal-time structure factor as entanglement witness based measures. In Sec. VII.1, we discuss the importance of realizing what is being experimentally accessed and how it affects the way one performs the disorder-average in their analysis. In particular, we highlight the importance of knowing how to perform the disorder-average keeping in mind the observables accessible in an experiment. We point that this is especially crucial when one is trying to rule out disorder induced phases by using certain entanglement measures. Following which, in Sec. VII.3, we find that the low-temperature growth of entanglement witness based measure turns out to be very useful to characterize these phases.
@@ VII.1 The order of disorder-average matters for concurrence
If one goes by our previous results for concurrence then one expects to see beyond nearest neighbor 2-partite entanglement in the RS phase. Does this imply that if experimentally we do not observe further nearest neighbor concurrence can we rule out the presence of a disorder induced phase? On the practical side, can the further nearest neighbor entanglement even be observed in INS? Unfortunately, the answer to both these questions turns out to be negative. Typically, in experimental situations such as in INS, we access disorder-averaged spin correlations. What this means is that when characterizing entanglement from spin correlation for example, one does not access the entanglement per disorder realization. While this distinction might feel academic, for measures like concurrence, the order of how one performs to disorder-average becomes important if one is comparing a theoretical model against the experimental predictions.
To begin with, let’s look at the finite temperature behavior of concurrence with the way we have been evaluating it as explained in Sec. VI.1. In Fig. 5, we find that up to some threshold temperatures the RS phase has finite pairwise entanglement from further neighbors whereas the TLL phase has only nearest neighbor pairwise contribution consistent with what we saw before.
@@
Figure 5: Demonstrating how the order of disorder-average affects the concurrence distribution in the random singlet (RS) phase. (a) Disorder-averaged concurrence observed for $1$st, $3$rd and $5$th neighbors as function of temperature for $N=20$. (b) Concurrence, obtained from disorder-averaged spin correlations, observed for the $1$st, $3$rd and $5$th neighbors as a function of temperature. For (a), concurrence is evaluated using Eq. (13) for every disorder realization and then averaged. On the other hand, for (b), concurrence is evaluated using Eq. (47). in which case the pairwise entanglement distribution of the RS phase mimics the Tomonga Luttinger liquid (TLL) phase i.e., becomes nearest neighbored. To help read the plots some of the error bars coming from disorder averages have been suppressed.
@@
where $\bar{g}^{zz}_{ij}$ is the disorder-averaged spin correlation function evaluated as per Eq. (35). From the experimental perspective, it is this concurrence that would be accessed. In Fig. 5, we perform this and find that RS has only nearest neighbor 2-partite entanglement just like the TLL. We do not observe any finite beyond nearest neighbor concurrence as one might have expected. What this means is that experimentally one would not be able to observe beyond nearest neighbor pairwise entanglement if they access disorder-averaged spin correlations. Thus ruling out the RS phase by looking at the 2-tangle or concurrence and observing no beyond nearest neighbor 2-partite entanglement is incorrect. We also observe in Fig. 5b that the nearest neighbor concurrence vanishes at the same temperature for both the TLL and RS phase.
Figure 6: Extracting central charge information from the low-temperature behavior of concurrence for the Tomonga Luttinger liquid (TLL) and random singlet (RS) phase respectively. The plots here are evaluated for $N=20$ spin chain. For the TLL phase, we make use of Eq. (48) and Eq. (49) to fit the observed low-temperature behavior. For the RS phase, we make use of Eq. (48) and Eq. (52) for the observed low-temperature behavior. Note that here we are evaluating the concurrence from disorder-averaged spin correlations [Eq. (47)] as one would access in experiments like inelastic Neutron scattering (INS). For ease of viewing, the error bars have been suppressed.
@@ VII.2 The low-temperature behavior of concurrence contains central charge
On the other hand, it has been shown that the RS phase is characterized by an effective central charge [98, 99, 94, 75]
@@
We conjecture that it is possible to access this information from the concurrence of the RS phase where we employ the disorder-average that is experimentally relevant. We find the following ansatz fits very well with the observed low-temperature behavior for the RS phase
@@ VII.3 The low-temperature growth of multi-partite entanglement is a useful probe
We saw that the zero temperature QFI, Fig. 4, predicts both the TLL and RS phase to be multi-partite entangled. Thus, stating whether a phase is $m$-partite or some other is $l$-partite is not a very useful probe if we want to distinguish them experimentally. Instead, we propose to investigate the low-temperature behavior of multi-partite entanglement witness measures. We employ the equal time structure factor for this task. As evident in Figs. 7, the growth of entanglement at low-temperatures is qualitatively quite different. For the TLL, the low-temperature scaling behavior is well known [100]
@@
On the other hand, we find the growth of entanglement to be quite slow in our disorder induced phase. We propose the following empirical behavior,
@@
Figure 7: Comparison of the low-temperature behavior of the equal time structure factor for the Tomonga Luttinger liquid (TLL) and random singlet (RS) phase. (a) Equal time structure factor of both the TLL and RS phases observed as function of $\log(\beta)$ for $N=16,20,24$ chains. (b) Testing the low-temperature behavior of equal time structure factor for the TLL against the analytic prediction Eq. (55). (c) Testing the low-temperature behavior of equal time structure factor for the RS against the ansatz Eq. (57) for $N=20$ spin chain. The fitting parameters are given in Eq. (56) and Eq. (58). The fits were checked within a fixed interval $1.90.09<\log(\beta)<$ 1.9 for both the TLL and RS phases. This choice was made owing to the finite size effects in the TLL phase for temperatures below the lower bound. For ease of viewing, the error bars have been suppressed.
@@ VIII Discussion
VIII.1 Multi-partite entanglement localization and understanding RS ‘phase’
Within the SDRG approach, the RS phase is pictured as a tensor product of singlets formed out of effective spins per disorder realization Fig. 8a [65]. On the other hand, both RT Fig. 3 and QFI Fig. 4, point to the RS phase to be multi-partite entangled. Is there an apparent contradiction here? No, there isn’t. The key here is to realize that the singlets in the SDRG picture is between effective spins [65]. The short range physics is scaled away in the renormalization procedure and it is this short length scale physics that contributes to our observation of multi-partite entanglement.
Figure 8: An illustration of the random singlet (RS) phase. (a) The schematic picture of the RS phase, as described within the strong disorder renormalization group (SDRG), where pairs of entangled effective spins (black) are connected by orange colored arcs. The green shaded ovals highlight that the effective spins are made up of multiple real space spins. (b) Schematic picture of the RS phase predicted from our results, where multi-partite clusters (mc) are represented by green shaded ovals of size $l_{mc}$ alongside pairwise entangled sites (blue) connected by orange colored arcs. These mc are localized to distances below the crossover length $L_{c}$.
The SDRG framework provides an extremely good description of the long-wavelength behavior of a disordered spin chain [101, 99]. However this long-wavelength picture, in which correlations are controlled by an infinite–disorder fixed point, should be understood to come into play above a crossover length, $L_{c}$ [101]. This crossover can be observed in Fig. 4a, as the momentum
@@ VIII.1 Multi-partite entanglement localization and understanding RS ‘phase’
below which the RS phase and the TLL are practically indistinguishable. What this means is that at distances shorter than $L_{c}$, spins can be entangled in clusters that contribute to the observed multipartite entanglement, as illustrated in Fig. 8b. However for distances $\gg L_{c}$, these ‘multi-partite’ clusters are completely suppressed, where now the dominant entanglement is at most 2-partite. At short length scales, the multi-partite entanglement plays a role, as witnessed by the QFI density [Fig. 4a]. We understand this as the localization of the multi-partite entanglement.
We can turn this observation into a bound on QFI density. If the RS phase, as understood within the SDRG, is at most 2-partite entangled, this places a limit on the entanglement contributed by long length-scales. We therefore expect
@@
A violation of this inequality would rule out the RS phase.
To summarize the above discussion: the RS phase contains multi-partite entanglement at short length scales. This is not in contradiction with the SDRG approach, where multi-partite entanglement information is absorbed within effective spin-1/2 moments, which go on to form singlets at long length scales. And as we have documented, the multi-partite entanglement found in numerical simulations of disordered spin chains originates in contributions below a crossover scale i.e., from short–range physics. In short, while multi-partite entanglement is not relevant to correlations at the longest length scales, which are controlled by the SDRG fixed point, it does exist. As a consequence, the observation of finite multi-partite entanglement using QFI cannot be used to rule out a RS phase.
VIII.2 Infinite disorder and Thermodynamic limit.
@@
Figure 9: Observing the decreasing entanglement depth with increasing $\delta$ in the random singlet (RS) phase. (a) Quantum Fisher information (QFI) density as function of q for $\delta=1.0,1.5,2.0,2.5$ in a $N=30$ spin chain. (b) Residual tangle (RT) as a function of $\frac{1}{\delta}$ for $N=30$ chain, where the line is used as a guide. While an increasing $\delta$ reduces the entanglement depth, for $\delta=2.5$ and $N=30$ spin chain the QFI density at the $q=\pi$ still detects a multi-partite entangled RS phase.
@@
where $C^{*}$ is some constant. For real materials, a finite $\delta$ would already imply an entanglement depth larger than two. For instance, in [102], it was reported that Ba5CuIr3O12 single crystals can characterize a RS phase with a $\delta\approx 2.5$.
VIII.3 Experimentally ruling out RS phase is a subtle business
@@
The RS phase is multi-partite entangled. This could have been inferred from the results in [94, 103], in hindsight. While the 2-partite entanglement of this phase may indeed be the relevant entanglement setting to understand the effects of disorder, it does also have a multi-partite contribution from length scales below the crossover length. Even though the multi-partite nature of the phase may not be relevant to understand the effects of disorder at long length scales, these contributions do not simply disappear either. Consequently, ruling out RS phase under the assumption that it has only 2-partite entanglement is incorrect. Furthermore, we showed how one might be misled into ruling out RS phase by the absence of any concurrence beyond the nearest neighbors. One might incorrectly conclude the absence of long range singlets from such an observation.
@@
Convex roof measures like concurrence depend quite a lot on the way the disorder-average is performed. This is due to the presence of the ‘$\max$’ function in the definition Eq. (10). The operation of averaging over disorder realizations does not commute with the max operation. Whether one takes the disorder-average inside the max function, or outside it, the end results are completely different. Why is this important? For quenched disordered systems, the spin correlations are already disorder-averaged. Thus if one were to access spin correlations from experiments like INS and from those tried evaluated concurrence using Eq. (47) then one would observe Fig. 5b as opposed to Fig. 5a. The former corresponds to the case when concurrence has been evaluated from disorder-averaged spin correlations and the latter is the one where we evaluate concurrence for each disorder realization and then average it. In the former case, the RS phase shows no further neighbor concurrence, essentially mimicking the TLL phase at the 2-partite level Fig. 2c. In [59], the authors incorrectly ruled out RS phase by observing vanishing concurrence and multi-partite entanglement using QFI. However, both of these observations cannot rule out the RS phase as shown in Sec. VI.2 and VII.1. This point will be discussed further in a follow-up work [104].
VIII.4 What can be useful experimentally?